alones (alones):

@mathstudent55 Check

alones (alones):

\[11x = 4 - 3x^2\] by factoring For which I got \[(\frac{ 1 }{ 3 },-4)\]

OpenStudy (mathstudent55):

\(11x = 4 - 3x^2\) \(3x^2 + 11x - 4 = 0\) \((3x - 1)(x + 4) = 0\) \(3x - 1 = 0\) or \(x + 4 = 0\) \(3x = 1\) or \(x = -4\) \(x = \dfrac{1}{3} \) or \(x = -4\)

OpenStudy (mathstudent55):

You are correct.

Nnesha (nnesha):

you can substitute x for each x value to check the answer if both sides are equal at the end then yes the answer is correct \[\large \rm 11(\frac{1}{3})= 4-3(\frac{1}{3})^2\]\[\large \rm 11(-4)= 4-3(-4)^2\]

alones (alones):

Oh okay thank you cx